How much will $100 grow to if invested at a continuously compounded interest rate of 7.75% for 9 years? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

How much will $100 grow to if invested at a continuously compounded interest rate of 9% for 7.75 years? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

HI,

As per policy we will solve only first question here

Present valueP =$100

Time t=9 years

Rate=7.75% continuous compounded

Future value F= P *e^(rt)

= 100*e^(9*0.0775)

=100*e^(0.6975)

=100*2.00872

=$200.87

Thanks